1.Our hydrocarbon has the molecule or formula C6H12 which of the following statements interpret this information correctly

A. 6 g of C combines with 12 grams of H to form this compound

B. One mol of this compound contains 6 mol of C and 12 mol of H

C. The ratio of C:H in this compound is 1:2

2.Quinine, which was 1st isolated in 1820 from the bark of a Cinchona tree, has been used for hundreds of years to treat malaria. An elemental analyst of Quinine shows that it contains 73.99% C ,7.50% H, 8.62% N, and 9.89% O by mass.

The 1st step in determining the empirical formula of this compound is to convert these masses to moles to do this you should____?

A. I seem that we have 100 g of the compound and treat the percentages as masses.

B. A scene that these percentages are equal the number of moles of each element present

C. Reduce those percentages to the lowest way she'll poll numbers.

3. The next step is to calculate the number of moles of each element present in 100.0 g of compound. giving each number of moles to 4 significant digits.

4. Now use the calculation from part free to determine the empirical formula of the compound and to the empyrical formula listing the elements in order CHNO.

5. For molar mass of quinine is 324.42 g/mol the correct molecular formula for quinine is...

C10H12NO
C20H24N2O2
C20H12NO2
C20H12NO
Answer #1

1.

Formula is C6H12.

Hence, mole ratio of Carbon and hydrogen in 1 mole of the compound is C :H = 6:12.

Hence, option B is correct statement.

2.

Percent mass is given. So total mass can be treated as 100

and percentages mass can be treated as masses.

Option A is correct.

3.

Element % mass atomic mass mole mole ratio
C 73.99 12

(73.99/12)

= 6.166

(6.166/0.6157)

= 10

H 7.50 1

(7.50/1)

= 7.500

(7.500/0.6157)

= 12

N 8.62 14

(8.62/14)

= 0.6157

(0.6157/0.6157)

= 1

O 9.89 16

(9.89/16)

= 0.6181

(0.6181/0.6157)

= 1

4.

Hence,

C: H : N: O

= 10:12:1:1

So, empirical formulais C10H12NO.

5.

Empirical formula weight

= 12 × 10 + 12× 1 + 14 + 16

= 162 g/mol.

Now, empirical formula weight × n = molecular weight

Or, 162 × n = 324.42

Or, n = (324.42/162) = 2.

Therefore molecular formula is

(C10H12NO)2 = C20H24N2O2 .